# Forecasting intraday foreign exchange volatility with functional GARCH approaches

### v2

@@ 5.1 Comparison of FGARCH-type models
First, we assess the predicability of FGARCH-type models. In total, we consider eight forecasting models that are the permutations of two model specifications. These are FGARCH(1,1)11(1,1)( 1 , 1 ) and the FGARCH-X model, along with four dimension reduction basis functions: functional principal components (FPCA), dynamic functional principal components (DFPCA) of Shang and Kearney (2022), long-range dependent functional principal component (LFPCA), and multi-level functional principal components (MFPCA). We demean the OCIDR curves to obtain $\widetilde{y}_{t}(u)$ by removing the functional mean. The parameters $K$ in (2.3) and $K^{c}$ and $K^{j}$ in (4.2), that is, the dimension of our basis functions, have to be determined when estimating the forecasting models. We select the appropriate number of these parameters using the eigenvalue ratio method (2.6) for each model estimation. This method indicates that using only the first functional principal component, that is, Basis function 1, is adequate in all cases. To inform our parameter selection, we decompose the data-driven basis functions from squared intraday returns. Table 5.1 shows that the first basis explains most of the total variation. The first basis of DFPCA roughly accounts for 80%, with the first basis generally explaining 95% for the other approaches. The results are produced using the first training sample (07-January-2014 to 17-May-2017), and the observations generalise when adopting other training samples. Besides, by testing the goodness-of-fit of the FGARCH(1,1)11(1,1)( 1 , 1 ) and FGARCH-X models based on the chosen first data-driven bases, we learn that the models are adequate in explaining the conditional heteroscedasticity of the FX OCIDR curves. The results of the goodness-of-fit test by Rice et al. (2020b) are omitted here to save space but can be obtained upon request.
@@
Figure 5.1 displays the intraday volatility forecasts of USD/EUR obtained from the FGARCH and FGARCH-X models using MFPCA bases, as well as the squared OCIDR curves. We find that both the FGARCH and FGARCH-X models forecast intraday volatility to be higher towards the end of the day. This is because the bases derived from the training sample capture the same pattern, with cumulative intraday returns generally increasing as the day progresses, which is also reflected in the squared OCIDR curves. Additionally, the intraday volatility forecasts are notably higher in 2018 and 2020, capturing features similar to those of the squared OCIDR curves. This increased volatility corresponds to multiple US interest rate hikes, the US-China trade war, political uncertainty in Europe in 2018, and the outbreak of the COVID-19 pandemic that led to significant central bank interventions in 2020. Besides, after accounting for the bid-ask spread, the FGARCH-X model forecasts volatility on a relatively lower scale, although the overall pattern remains similar to that predicted by the FGARCH model.
Figure 5.2 plots the mean and median volatility forecasts of eight candidate models and the squared OCIDR curves over 24-hour trading sessions to enhance our understanding of the distribution of volatility forecasts. We find that all forecasts reasonably attempt to represent the dynamics of squared OCIDR curves. Remarkably, the MFPCA-embedded models capture spikes at the end of the trading day, followed by the LFPCA-associated models that show slightly better capabilities than the TFPCA and DFPCA-embedded models. Comparatively, FGARCH-X models provide forecasts closer to the squared OCIDR curves.
Figure 5.2: Plots of pointwise Mean and Median of intraday volatility forecasts and squared OCIDR curves of USD-EUR over 24-hour trading sessions obtained from eight FGARCH and FGARCH-X models with the basis functions are derived by using the TFPCA, DFPCA, LFPCA, and MFPCA approaches. Our first training sample of 07-January-2014 to 17-May-2017 and a forecasting sample of 18-May-2017 to 30-September-2020 are specified with a rolling forecast window dynamically updated weekly. The left sub-figure displays the plots of mean curves, and the right sub-figure shows the plots of median curves.
@@
Table 5.2 shows the loss functions based on our forecasting errors. The comparative performances are mostly consistent when measured using either MSFE or QLIKE loss functions. Notably, in both Panels A and B, the models embedded with MFPCA bases (which account for cross-dependence among FX rates) generally produce the smallest forecasting errors for USD/EUR and USD/JPY. This evidences that our multivariate approach disentangles common variation from currency-specific information, capturing joint shocks for volatility modelling more effectively. The results for USD/GBP are somewhat mixed, with the smaller MSFE suggesting using TFPCA or LFPCA bases. We also notice that the MFPCA bases are more effective in improving the accuracy of the FGARCH(1,1) model when compared with the FGARCH-X model. This may be because the cross-dependence variations driven by the MFPCA bases do not properly consider the variations in the covariate X. Moreover, the models using an LFPCA basis, particularly in Panel B, outperform the TFPCA and DFPCA bases models, demonstrating the benefit of explaining the long-range dependence of the OCIDR curves. The models relying on a DFPCA basis perform slightly better than the TFPCA counterparts in forecasting intraday and inter-daily conditional volatility. Furthermore, the MSFE and QLIKE generally give consistent results when assessing the inter-daily conditional volatility forecasts. The results are more mixed in USD/GBP intraday conditional volatility forecasts, where some inconsistency arises from specifying the TFPCA or LFPCA bases. Lastly, to compare the FGARCH and FGARCH-X models, we observe that the FGARCH-X models generate considerably lower forecasting errors than the FGARCH model in both the intraday and inter-daily conditional volatility forecasts. This is because different FPCA methods serve as data-driven approaches that only capture the dynamics of the squared process of OCIDR. By incorporating bid-ask spread information, the FGARCH-X model offers the opportunity to correct the level of conditional volatility modelling, enhancing its predictive accuracy.
We then apply the Model Confidence Set (MCS) of Hansen et al. (2011) and the Diebold and Mariano’s (1995) test to assess the relative performances in multi-model and pairwise models, respectively. Table 5.3 shows the forecasting models designated to lie in the superior set of models at 95% significance level for each FX rate. Our findings indicate that the intraday bid-ask spread information significantly improves predictability by reducing forecasting errors. The FGARCH-X model is always elected to be the superior model in forecasting either intraday conditional volatility in Panel A or inter-daily conditional volatility in Panel B.
Table 5.3: Results of model confidence set tests (Hansen et al., 2011) based on the MSFE and QLIKE forecasting errors presented in Equations (5.2) and (5.1) at 95% confidence set level. FGARCH labelled FG-, and FGARCH-X, labelled X-, models are specified with four types of data-driven basis functions: truncated functional principal components (TFPCA), dynamic functional principal components (DFPCA), long-range dependent functional principal components (LFPCA), and multi-level functional principal components (MFPCA). Our first training sample of 07-January-2014 to 17-May-2017 and a forecasting sample of 18-May-2017 to 30-September-2020 are specified with a rolling forecast window dynamically updated weekly. Panel A displays the models falling into the confidence sets that forecast intraday conditional volatility. Panel B shows the models that fall into the confidence sets that forecast inter-daily conditional volatility.
@@
Our findings complement the literature and provide further evidence that intraday bid-ask spread information provides predictability to conditional volatility in FX markets. In addition, to compare different basis functions used, the models associated with the MFPCA approach are still chosen most often for USD/EUR and USD/JPY. For the USD/GBP rate, the FGARCH-X model using the LFPCA basis mostly falls into all confidence sets, with the exception that the TFPCA is selected when the QLIKE measurements are used, which follows the results of Table 5.2. Furthermore, it is interesting to note that, according to Table 3.1, USD/GBP is more long-range dependent compared with the other two rates at the second moment with the long-memory parameter $\widehat{a}_{x^{2}}=0.33$, which could be the reason that the models associated with the LFPCA basis outperform for USD/GBP.
@@
Table 5.4 shows the results of the Diebold-Mariano test on pairwise forecasting models. Twelve model pairs are compared, including ten pairs of the FGARCH and FGARCH-X models, with TFPCA and DFPCA serving as benchmarks. Additionally, two pairs of FGARCH and FGARCH-X models are compared, where the same types of LFPCA and MFPCA bases are applied. In terms of intraday conditional volatility forecasts, the results demonstrate that, in a few cases for all FX rates, the benchmark models significantly underperform the model, relying on the other proposed bases. The MFPCA, as the superior basis, consistently predicts more accurate conditional volatility for FGARCH models in USD/EUR. However, the relative performance of the FGARCH-X models does not significantly differ from that of models using other bases. Finally, to compare FGARCH and FGARCH-X models, the statistics overall indicate a positive sign towards the FGARCH-X models, although they are rarely significant at the 95% level in USD/GBP and USD/JPY. These patterns are broadly retained for the inter-daily conditional volatility forecasts.
@@ 5.2 Comparison with realised volatility-based models
Next, to evaluate the benefits of incorporating intraday trading movements for inter-daily volatility forecasts, we compare the proposed FGARCH-type models with nine traditional volatility models, including the GARCH(1,1), GJR-GARCH(1,1), GARCH-X, Fractionally Integrated-GARCH (FIGARCH), Heterogeneous Autoregressive (HAR) (Corsi, 2009), HAR-X, HAR-Autoregressive Fractionally Integrated Moving Average(ARFIMA), Realised GARCH (RGARCH) (Hansen et al., 2012), and the GARCH-MIDAS (Engle et al., 2008, 2013). Note that these comparisons only apply to inter-daily volatility forecasts.
@@
[…]
where $\mbox{mid}_{t}$ is the average of bid and ask prices at day $t$. The fractional integration order $d$ in FIGARCH(1,d,1) and HAR-ARFIMA(1,d,1) are optimally selected using the functions from rugarch and arfima in the statistical software [Uncaptioned image]. Also, in realised volatility-based models, we obtain realised volatility by summing squared intraday log returns with the squared overnight return, similar to how it was used in the main manuscript. Following Engle et al. (2013), we use monthly realised volatility for the GARCH-MIDAS model to capture long-term variations. The realised volatility-based models are mainly implemented using functions from the packages HARModel (Sjoerup, 2019), rugarch (Galanos, 2024) and rumidas (Candila, 2024) in the statistical software [Uncaptioned image] (R Core Team, 2024). Similar to Section 5 of the main manuscript, we use the entire sample from 07-January-2014 to 30-September-20 and train each model candidate with a rolling window of 600 observations, resulting in the out-of-sample period being from 18-May-2017 to 30-September-2022. Similarly, each model is re-estimated every week to adapt to market changes. From the eight FGARCH-type models, we select the FGARCH-MFPCA model and the FGARCH-X models using LFPCA and MFPCA bases as representative models for this comparison.
Table 5.5 exhibits the MSFE and QLIKE forecasting errors of eleven model candidates, which are calculated by using (5.1) from the main context with the realised volatility estimating true volatility. Except for the HAR-X model producing the smallest MFSE error for USD/GBP, the overall results indicate that the FGARCH-X model outperforms other model candidates. The realised volatility-based models generally outperform scalar GARCH models, which also beat the FGARCH model. The FGARCH model shows comparable performances to scalar GARCH models for USD/GBP and USD/JPY, but the errors get off-scale for USD/EUR. Besides, we observe that, aside from the FGARCH-X models, the HAR-X model outperforms in producing smaller forecasting errors, further supporting the predictability of bid-ask spreads to FX volatility forecasting. Comparatively, by incorporating all the intraday trading information, the FGARCH-X shows further superiority. Figure 5.3 displays the plots of realised volatility and inter-daily volatility forecasts for nine model candidates, along with the FGARCH-X-MFPCA, across three FX pairs. The plotting patterns are consistent with the findings revealed from Table 5.5 that the FGARCH-X volatility forecasts are closer to the realised volatility.
@@ 6 A VaR corrected intraday trading strategy
where $\widehat{\sigma}_{t+1}(u)$ is the intraday volatility forecasts obtained from the models, and $\widehat{\varepsilon}^{\zeta}(u)$ is the $\zeta$th unconditional quantile of the error process. Rice et al. (2020b) suggest three types of error processes for $\widehat{\varepsilon}^{\zeta}(u)$, including a Gaussian process, generalised Pareto distributed process, and the empirical process obtained by bootstrapping the fitted error curves. Here, we estimate $\widehat{\varepsilon}^{\zeta}(u)$ through bootstrapping the empirical process with 10,000 replications.
@@
where $\mathbb{I}\{\cdot\}$ is a binary indicator function so that the process $Z_{t}^{\zeta}(u)$ is composed of binary values, with one standing for pointwise exceeding the intraday interval. The intraday VaR curves are valid only if the process $Z_{t}^{\zeta}(u)$ is unbiased from the nominal level $\zeta$ and independent along $t$. To evaluate the validity of intraday VaR forecasts as well as to guide long and short trading strategies, we test the unbiasedness and independence hypotheses when VaR curves are computed at nominal quantiles $\zeta=0.01$ and $0.99$ for potential intraday risk management tools.
The left part of Table 6.1 documents the backtest results for the intraday VaR forecasts at nominal quantiles $\zeta=0.01$. Nearly all model candidates using the proposed bases produce unbiased intraday VaR curves, and their violation processes cannot reject the independence hypothesis at lags $H=1$, $5$, $10$ and $20$. The only exception occurs for the asset USD/JPY where the intraday VaR associated with the LFPCA basis function cannot pass the unbiasedness test. Overall, the results indicate valid intraday VaR forecasts, which can be used as risk management tools for intraday trading strategies. We use the intraday VaR of $\zeta=0.01$ and $\zeta=0.99$ to control for tail risks in long and short trading strategies. These findings generally remain consistent for $\zeta=0.99$, so we skip reporting similar results.
Table 6.1: $P$-values of the unbiasedness and independence backtests for intraday VaR curves with nominal quantile $\zeta=0.01$ and $0.0025$. FGARCH and FGARCH-X models are specified with the proposed data-driven basis functions: long-range dependent functional principal components (LFPCA) and multi-level functional principal components (MFPCA). Our first training sample of 07-January-2014 to 17-May-2017 and a forecasting sample of 18-May-2017 to 30-September-2020 are specified with a rolling forecast window that is dynamically updated weekly. The lag lengths of the independence test are set at $H=1$, $5$, $10$ and $20$, with bold values being significant at the 5% significance level.
@@
To further investigate the effect of extreme quantiles on the intraday VaR, the right part of Table 6.1 displays the backtesting results for $\zeta=0.0025$. The validity of our intraday VaR forecasts generally remains for USD/EUR and USD/JPY, but it deteriorates for USD/GBP. Most forecasting models cannot pass the unbiasedness backtesting, although the corresponding violation processes exhibit the desirable independent property. Despite this, the findings of $\zeta=0.0025$ are roughly consistent with those presented for $\zeta=0.01$.
We now discuss an intraday asset allocation problem that benefits from constructing valid intraday VaR curves. Here, we aim to examine whether intraday VaR provides a sufficient risk management buffer to prevent large economic losses inherent in intraday trading strategies. Thus, we introduce a simple VaR-corrected intraday trading strategy. The principle is that for each trading day, we open a position at the minimum point of the predicted OCIDR curve and close the position at the maximum point of the predicted OCIDR curve before the close of trading. When considering a long trading strategy, the asset is bought at the predicted minimum point and sold at the maximum point, and vice versa for a short trading strategy. The strategy is only executed when the range of $\widehat{y}_{t+1}(u)$ exceeds its mean. Otherwise, active trading is suppressed, and the return is zero on that trading day.
@@
In further unreported analysis, we calculate historical functional means with lagged periods of a quarter (51 days), half-year (102 days), one year (204 days) and two years (408 days) based on four trading days a week. The results indicate improved performance when the half-year look-back period is deployed.
Next, we incorporate the VaR forecasts to mitigate risk during trading. Inspired by the idea of interval-valued information set, we only trade on the day that meets the condition $|\max{\widehat{y}_{t+1}(u)}-\min{\widehat{y}_{t+1}(u)}|\geq\frac{1}{J}\sum_{j=1}^{J}\widehat{y}_{t+1}(u_{j})$. On each active trading day with open trading positions, if the OCIDR realisation hits our intraday VaR forecasts, we close the position when the violation occurs. To mimic the actual trading in practice, we apply the usual FX market transaction cost of $0.0003\%$ in the simulation. The entire strategy, therefore, is as detailed in Algorithm 1 below. The upper subplot of Figure 6.1 shows examples of a long trading day, illustrating the scenarios that trigger VaR corrections on the benchmark trading strategy.
@@
The return curves are presented as a percentage by multiplying by 100 as shown in (3.1). Adopting a more practitioner-focused convention, here we consider the actual value rather than the percentage form to calculate the trading performance statistics. Table 6.2 presents the results for the benchmark and FGARCH-VaR-corrected trading strategies using MFPCA bases for USD/EUR, USD/GBP, and USD/JPY. We see improvements in annual returns, Sharpe ratios, and reduced maximum drawdown values across all currencies and models. Notably, the trading strategies for USD/GBP and USD/JPY produce positive returns after intraday VaR corrections. In Panel A of the long trading strategy, the FGARCH-X model shows superiority across most scenarios, which is consistent with the findings of smaller forecasting errors, as discussed in Section 5. The results in Panel B of the short trading strategy show slightly mixed preferences between FGARCH and FGARCH-X, but the general pattern aligns with the long trading performance, indicating the outperformance of VaR-corrected trading strategies.
@@
To enhance our understanding, we also plot the cumulative returns of the long trading strategy for USD/GBP as an example. The lower subplot of Figure 6.1 shows the comparisons between the benchmark and VaR-corrected strategies. We find that the three holding strategies generally follow a similar pattern. This is because the VaR-corrected strategy, as a risk management tool, only activates when there is a violation of the 1% intraday VaR curve. Otherwise, the holding returns are the same as the benchmark. The VaR-corrected strategy yields higher returns because it avoids sharp declines during extreme market events. Notably, during the outbreak of COVID-19, the benchmark experienced a dramatic decline in returns, whereas the VaR-corrected strategy successfully avoided this drop. Overall, the results show that we can provide a useful intraday risk management tool for trading activities by accurately forecasting intraday volatility.

### v3

@@ 4.1 Comparison of FGARCH-type models
First, we assess the predicability of FGARCH-type models. In total, we consider eight forecasting models that are the permutations of two model specifications. These are FGARCH(1,1)(1,1) and the FGARCH-X model, along with four dimension reduction basis functions: functional principal components (FPCA), dynamic functional principal components (DFPCA) of Shang and Kearney, (2022), long-range dependent functional principal component (LFPCA), and multi-level functional principal components (MFPCA). We demean the OCIDR curves to obtain $\widetilde{y}_{t}(u)$ by removing the functional mean. The parameters $K$ in (3.3) and $K^{c}$ and $K^{j}$ in (3.8), that is, the dimension of our basis functions, have to be determined when estimating the forecasting models. We select the appropriate number of these parameters using the eigenvalue ratio method (3.6) for each model estimation. This method indicates that using only the first functional principal component, that is, Basis function 1, is adequate in all cases. To inform our parameter selection, we decompose the data-driven basis functions from squared intraday returns. Table 4.1 shows that the first basis explains most of the total variation. The first basis of DFPCA roughly accounts for 80%, with the first basis generally explaining 95% for the other approaches. The results are produced using the first training sample (07-January-2014 to 17-May-2017), and the observations generalise when adopting other training samples. In addition, by testing the goodness-of-fit of the FGARCH(1,1)(1,1) and FGARCH-X models based on the chosen data-driven bases, we learn that the models are adequate to explain the conditional heteroskedasticity of the FX OCIDR curves. The results of the goodness-of-fit test by Rice et al., 2020b are omitted here to save space, but can be obtained upon request.
@@
Figure 4.1 displays the intraday volatility forecasts of USD/EUR obtained from the FGARCH and FGARCH-X models using MFPCA bases, as well as the squared OCIDR curves. We find that both the FGARCH and FGARCH-X models reasonably attempt to represent the dynamics of squared OCIDR curves, forecasting intraday volatility to increase toward the end of the trading day. This is because the bases derived from the training sample capture the same pattern, with cumulative intraday returns generally increasing as the day progresses, which is also reflected in the squared OCIDR curves. MFPCA-embedded model captures more spikes at the end of the trading day, the FGARCH-X models in average provide forecasts closer to the squared OCIDR curves. Additionally, intraday volatility forecasts are notably higher in 2018 and 2020, capturing features similar to those of squared OCIDR curves. This increased volatility corresponds to multiple US interest rate hikes, the US-China trade war, political uncertainty in Europe in 2018, and the outbreak of the COVID-19 pandemic that led to significant central bank interventions in 2020. In addition, after accounting for the bid-ask spread, the FGARCH-X model forecasts volatility on a relatively lower scale, although the overall pattern remains similar to that predicted by the FGARCH model.
@@
Table 4.2 shows the loss functions based on our forecasting errors. The comparative performances are mostly consistent when measured using either MSFE or QLIKE loss functions. Notably, in both Panels A and B, the models embedded with MFPCA bases (which account for cross-dependence among FX rates) generally produce the smallest forecasting errors for USD/EUR and USD/JPY. This evidences that our multivariate approach disentangles common variation from currency-specific information, capturing joint shocks for volatility modelling more effectively. The results for USD/GBP are somewhat mixed, with the smaller MSFE suggesting the use of TFPCA or LFPCA bases. We also notice that MFPCA bases are more effective in improving the accuracy of the FGARCH(1,1) model compared to the FGARCH-X model. This may be because the cross-dependence variations driven by the MFPCA bases do not properly consider the variations in the covariate X. Moreover, the models using an LFPCA basis, particularly in Panel B, outperform the TFPCA and DFPCA bases models, demonstrating the benefit of explaining the long-range dependence of the OCIDR curves. The models relying on a DFPCA basis perform slightly better than the TFPCA counterparts in forecasting intraday and inter-daily conditional volatility. Furthermore, the MSFE and QLIKE generally give consistent results when assessing the inter-daily conditional volatility forecasts. The results are more mixed in the USD/GBP intraday conditional volatility forecasts, where some inconsistency arises from specifying the TFPCA or LFPCA bases. Lastly, to compare the FGARCH and FGARCH-X models, we observe that the FGARCH-X models generate considerably lower forecasting errors than the FGARCH model in both the intraday and inter-daily conditional volatility forecasts. This is because different FPCA methods serve as data-driven approaches that capture only the dynamics of the squared OCIDR process. By incorporating bid-ask spread information, the FGARCH-X model offers the opportunity to correct the level of conditional volatility modelling, enhancing its predictive accuracy.
@@
We then apply the Model Confidence Set (MCS) of Hansen et al., (2011) and the Diebold and Mariano,’s (1995) test to assess the relative performances in multi-model and pairwise models, respectively. Table 4.3 shows the forecasting models designated to lie in the superior set of models at 95% significance level for each FX rate. Our findings indicate that the intraday bid-ask spread information significantly improves predictability by reducing forecasting errors. The FGARCH-X model is always elected to be the superior model to forecast either intraday conditional volatility in Panel A or inter-daily conditional volatility in Panel B.
Table 4.3: Results of model confidence set tests (Hansen et al.,, 2011) based on the MSFE and QLIKE forecasting errors presented in Equations (4.2) and (4.1) at 95% confidence set level. FGARCH labelled FG-, and FGARCH-X, labelled X-, models are specified with four types of data-driven basis functions: truncated functional principal components (TFPCA), dynamic functional principal components (DFPCA), long-range dependent functional principal components (LFPCA), and multi-level functional principal components (MFPCA). Our first training sample of 07-January-2014 to 17-May-2017 and a forecasting sample of 18-May-2017 to 30-September-2020 are specified with a rolling forecast window dynamically updated weekly. Panel A displays the models falling into the confidence sets that forecast intraday conditional volatility. Panel B shows the models that fall into the confidence sets that forecast inter-daily conditional volatility.
@@
Our findings complement the literature and provide further evidence that intraday bid-ask spread information provides predictability to conditional volatility in FX markets. In addition, to compare the different basis functions used, the models associated with the MFPCA approach are still chosen most often for USD/EUR and USD/JPY. For the USD/GBP rate, the FGARCH-X model using the LFPCA basis mostly falls into all confidence sets, with the exception that the TFPCA is selected when the QLIKE measurements are used, which follows the results of Table 4.2. Furthermore, it is interesting to note that, according to Table 2.1, USD/GBP is more long-range dependent compared to the other two rates at the second moment with the long-memory parameter $\widehat{a}_{x^{2}}=0.33$, which could be the reason why the models associated with the LFPCA basis outperform for USD/GBP.
Table 4.4 shows the results of the Diebold-Mariano test on pairwise forecasting models. Twelve model pairs are compared, including ten pairs of the FGARCH and FGARCH-X models, with TFPCA and DFPCA serving as benchmarks. Additionally, two pairs of FGARCH and FGARCH-X models are compared, where the same types of LFPCA and MFPCA bases are applied. In terms of intraday conditional volatility forecasts, the results demonstrate that, in a few cases for all FX rates, the benchmark models significantly underperform the model, relying on the other proposed bases. The MFPCA, as the superior basis, consistently predicts more accurate conditional volatility for FGARCH models in USD/EUR. However, the relative performance of the FGARCH-X models does not differ significantly from that of the models using other bases. Finally, to compare the FGARCH and FGARCH-X models, the statistics overall indicate a positive sign toward the FGARCH-X models, although they are rarely significant at the 95% level in USD/GBP and USD/JPY. These patterns are broadly retained for the inter-daily conditional volatility forecasts.
@@ 4.2 Comparison with realised volatility-based models
Next, to evaluate the benefits of incorporating intraday trading movements for inter-daily volatility forecasts, we also compare the proposed FGARCH-type models with nine traditional volatility models, including the GARCH(1,1), GJR-GARCH(1,1), GARCH-X, Fractionally Integrated-GARCH (FIGARCH), Heterogeneous Autoregressive (HAR) (Corsi,, 2009), HAR-X, HAR-Autoregressive Fractionally Integrated Moving Average(ARFIMA), Realised GARCH (RGARCH) (Hansen et al.,, 2012), and the GARCH-MIDAS (Engle et al.,, 2013). Note that these comparisons only apply to inter-daily volatility forecasts.
@@
$\mbox{Bid-ask Spread}_{t}=\frac{\mbox{ask}_{t}-\mbox{bid}_{t}}{\mbox{mid}_{t}},$
where $\mbox{mid}_{t}$ is the average of the bid and ask prices on day $t$. The fractional integration order $d$ in FIGARCH(1,d,1) and HAR-ARFIMA(1,d,1) are optimally selected using the functions from rugarch and arfima in the statistical software [Uncaptioned image]. Also, in realised volatility-based models, we obtain realised volatility by summing squared intraday log returns with the squared overnight return, similar to how it was used in the main manuscript. Following Engle et al., (2013), we use monthly realised volatility for the GARCH-MIDAS model to capture long-term variations. The realised volatility-based models are mainly implemented using functions from the packages HARModel (Sjoerup,, 2019), rugarch (Galanos,, 2025) and rumidas (Candila,, 2025) in statistical software [Uncaptioned image] (R Core Team,, 2025). Similar to Section 5 of the main manuscript, we use the entire sample from 07-January-2014 to 30-September-20 and train each model candidate with a rolling window of 600 observations, resulting in the out-of-sample period being from 18-May-2017 to 30-September-2022. Similarly, each model is re-estimated every week to adapt to market changes. From the eight FGARCH-type models, we select the FGARCH-MFPCA model and the FGARCH-X models using the LFPCA and MFPCA bases as representative models for this comparison.
Table 4.5 exhibits the MSFE and QLIKE forecasting errors of eleven model candidates, which are calculated by using (4.1) from the main context with the realised volatility estimating the true volatility. Except for the HAR-X model that produces the smallest MFSE error for USD/GBP, the overall results indicate that the FGARCH-X model outperforms other model candidates. The realised volatility-based models generally outperform scalar GARCH models, which also beat the FGARCH model. The FGARCH model shows comparable performances to the scalar GARCH models for USD/GBP and USD/JPY, but the errors are off-scale for USD/EUR. In addition, we observe that, apart from the FGARCH-X models, the HAR-X model outperforms in producing smaller forecasting errors, further supporting the predictability of bid-ask spreads to FX volatility forecasting. Comparatively, by incorporating all the intraday trading information, the FGARCH-X shows further superiority. Figure 4.2 displays the plots of realised volatility and inter-daily volatility forecasts for nine model candidates, along with the FGARCH-X-MFPCA, across three FX pairs. The plot patterns are consistent with the findings revealed in Table 4.5 that the FGARCH-X volatility forecasts are closer to the realised volatility.
@@ 5 A VaR corrected intraday trading strategy
[…]
@@
where $\mathbb{I}\{\cdot\}$ is a binary indicator function so that the process $Z_{t}^{\zeta}(u)$ is composed of binary values, with one standing for pointwise exceeding the intraday interval. The intraday VaR curves are valid only if the process $Z_{t}^{\zeta}(u)$ is unbiased from the nominal level $\zeta$ and independent of $t$. To evaluate the validity of intraday VaR forecasts as well as to guide long and short trading strategies, we test the unbiasedness and independence hypotheses when VaR curves are computed at nominal quantiles $\zeta=0.01$ and $0.99$ for potential intraday risk management tools.
The left part of Table 5.1 documents the backtest results for intraday VaR forecasts at nominal quantiles $\zeta=0.01$. Almost all model candidates using the proposed bases produce unbiased intraday VaR curves, and their violation processes cannot reject the independence hypothesis at lags $H=1$, 55, 1010 and $20$. The only exception occurs for the asset USD/JPY where the intraday VaR associated with the LFPCA basis function cannot pass the unbiasedness test. Overall, the results indicate valid intraday VaR forecasts, which can be used as risk management tools for intraday trading strategies. We use the intraday VaR of $\zeta=0.01$ and $\zeta=0.99$ to control for tail risks in long-and short-trading strategies. These findings generally remain consistent for $\zeta=0.99$, so we skip reporting similar results.
To further investigate the effect of extreme quantiles on intraday VaR, the right part of Table 5.1 displays the backtest results for $\zeta=0.0025$. The validity of our intraday VaR forecasts generally remains for USD/EUR and USD/JPY, but it deteriorates for USD/GBP. Most forecasting models cannot pass the unbiasedness backtest, although the corresponding violation processes exhibit the desirable independent property. Despite this, the findings of $\zeta=0.0025$ are roughly consistent with those presented for $\zeta=0.01$.
Table 5.1: $P$-values of the unbiasedness and independence backtests for intraday VaR curves with nominal quantile $\zeta=0.01$ and $0.0025$. FGARCH and FGARCH-X models are specified with the proposed data-driven basis functions: long-range dependent functional principal components (LFPCA) and multi-level functional principal components (MFPCA). Our first training sample of 07-January-2014 to 17-May-2017 and a forecasting sample of 18-May-2017 to 30-September-2020 are specified with a rolling forecast window that is dynamically updated weekly. The lag lengths of the independence test are set at $H=1$, 55, 1010 and 2020, with bold values being significant at the 5% significance level.
@@
We now discuss an intraday asset allocation problem that benefits from constructing valid intraday VaR curves. Here, we aim to examine whether intraday VaR provides a sufficient risk management buffer to prevent the large economic losses inherent in intraday trading strategies. Thus, we introduce a simple VaR-corrected intraday trading strategy. The principle is that for each trading day, we open a position at the minimum point of the predicted OCIDR curve and close the position at the maximum point of the predicted OCIDR curve before the close of trading. When considering a long-trading strategy, the asset is bought at the predicted minimum point and sold at the maximum point, and vice versa for a short-trading strategy. The strategy is only executed when the range of $\widehat{y}_{t+1}(u)$ exceeds its mean. Otherwise, active trading is suppressed and the return is zero on that trading day.
@@
In further unreported analysis, we calculate historical functional means with lagged periods of a quarter (51 days), half-year (102 days), one year (204 days), and two years (408 days) based on four trading days a week. The results indicate improved performance when the half-year look-back period is deployed.
Next, we incorporate the VaR forecasts to mitigate risk during trading. Inspired by the idea of interval-valued information set, we only trade on the day that meets the condition $\left|\max{\widehat{y}_{t+1}(u)}-\min{\widehat{y}_{t+1}(u)}\right|\geq\frac{1}{J}\sum_{j=1}^{J}\widehat{y}_{t+1}(u_{j})$. On each active trading day with open trading positions, if the OCIDR realisation hits our intraday VaR forecasts, we close the position when the violation occurs. To mimic actual trading in practice, we apply the usual FX market transaction cost of $0.0003\%$ in the simulation. The entire strategy, therefore, is as detailed in Algorithm 1 below. The upper subplot of Figure 5.1 shows examples of a long trading day, illustrating the scenarios that trigger VaR corrections on the benchmark trading strategy.
@@
Return curves are presented as a percentage by multiplying by 100 as shown in (2.1). Adopting a more practitioner-focused convention, here we consider the actual value rather than the percentage form to calculate the trading performance statistics. Table 5.2 presents the results for the benchmark and FGARCH-VaR-corrected trading strategies using MFPCA bases for USD/EUR, USD/GBP, and USD/JPY. We see improvements in annual returns, Sharpe ratios, and reduced maximum drawdown values across all currencies and models. Notably, the trading strategies for USD/GBP and USD/JPY produce positive returns after intraday VaR corrections. In Panel A of the long trading strategy, the FGARCH-X model shows superiority in most scenarios, which is consistent with the findings of smaller forecasting errors, as discussed in Section 4. The results in Panel B of the short trading strategy show slightly mixed preferences between FGARCH and FGARCH-X, but the general pattern aligns with the long trading performance, indicating the outperformance of VaR-corrected trading strategies.
@@
To enhance our understanding, we also plot the cumulative returns of the long trading strategy for USD/GBP as an example. The lower subplot of Figure 5.1 shows the comparisons between the benchmark and VaR-corrected strategies. We find that the three holding strategies generally follow a similar pattern. This is because the VaR-corrected strategy, as a risk management tool, only activates when there is a violation of the 1% intraday VaR curve, i.e., during bad days. Otherwise, the holding returns are the same as the benchmark. The VaR-corrected strategy yields higher returns because it avoids sharp declines during extreme market events. Notably, during the outbreak of COVID-19, the benchmark experienced a dramatic decline in returns, whereas the VaR-corrected strategy successfully avoided this drop. Overall, the results show that we can provide a useful intraday risk management tool for trading activities by accurately forecasting intraday volatility.
The implementation of the methodology applied in this article is available in the online repository: https://github.com/yzhao7322/FX-Multi-FGARCH.